OpenStudy (anonymous):

Find all of the solutions of the given equation in the interval (0,2π). sin(x-3π/4)+sin(x+3π/4)=√(2)

OpenStudy (anonymous):

Nevermind I figured it out:)

OpenStudy (anonymous):

would you know how to solve sin(x)=1/2 for \[0\le x \le \pi \div2\]?

OpenStudy (anonymous):

Uhm... no sory I actually ended up guessing for this question... Sorry wish I could help you:(

OpenStudy (anonymous):

*sorry

OpenStudy (anonymous):

np